# Imaging Shapes of Atomic Nuclei in High-Energy Nuclear Collisions

### v1

@@ Abstract
Atomic nuclei are self-organized, many-body quantum systems bound by strong nuclear forces within femtometer-scale space. These complex systems manifest a diverse set of shapes Bohr and Mottelson (1998); Heyde and Wood (2016), traditionally explored via non-invasive spectroscopic techniques at low energies Cline (1986); Yang et al. (2023). Their instantaneous shapes, obscured by long-timescale quantum fluctuations, are considered not directly observable at low energy. We introduce a complementary method, collective flow assisted nuclear shape imaging, to image the nuclear global shape by colliding them at ultrarelativistic speeds and analyzing the collective response of outgoing debris. This technique captures a collision-specific snapshot of the spatial matter distribution in the nuclei, which, through the hydrodynamic expansion, leaves imprints on the particle momentum distribution patterns observed in detectors Shuryak (2017); Busza et al. (2018). We benchmark this method in collisions of ground state Uranium-238 nuclei, known for its elongated, axial-symmetric shape. Our findings, while confirming an overall deformation broadly consistent with prior low-energy experiments, also indicate a small deviation from axial symmetry in the nuclear ground state. This approach marks a new way of imaging nuclei, especially those with uncertain shape characteristics, and refines initial conditions in high-energy nuclear collisions. It tackles the important issue of nuclear structure evolution across various energy scales.
More than 99.9% of the visible matter in the cosmos resides in the center of atoms – the atomic nuclei composed of nucleons (protons and neutrons). Our present knowledge of their global structure primarily comes from a variety of spectroscopic or scattering experiments Cline (1986); Yang et al. (2023); Hofstadter (1956) at beam energies below hundreds of MeV per nucleon. These studies reveal that most nuclei are ellipsoidally deformed, with greater deformation in nuclei distant from magic numbers (2, 8, 20, 28, 50, 82, and 126) Möller et al. (2016). Investigating this deformation across the nuclear chart has been an important area of research over many decades, and is crucial for topics such as nucleosynthesis Schatz et al. (1998), nuclear fission Schunck and Regnier (2022), and neutrinoless double beta decay ($0\nu\beta\beta$Engel and Menéndez (2017).
Within a collective model picture, the ellipsoidal shape of a nucleus with mass number $A$ is defined in the intrinsic (body-fixed) frame, where its surface $R(\theta,\phi)$ is described by Bohr and Mottelson (1998),
@@
Here, $R_{0}\approx 1.2A^{1/3}$ fm represents the nuclear radius (about 5–8 fm for large nuclei). The spherical harmonics in the real basis $Y_{l,m}(\theta,\phi)$, the quadrupole deformation magnitude $\beta_{2}$, and the triaxiality parameter $\gamma$ define the nuclear shape. The $\gamma$ parameter, spanning $0^{\circ}$ to $60^{\circ}$, controls the ratios of principal radii. Specifically, $\gamma=0^{\circ}$ corresponds to a prolate shape, $\gamma=60^{\circ}$ an oblate shape, and values in between $0^{\circ}<\gamma<60^{\circ}$ to a triaxial shape. While most nuclei are axially-symmetric (prolate or oblate) or have a fluctuating $\gamma$ value ($\gamma$-soft), rigid triaxial shape is uncommon Toh et al. (2013). An example of an axial-symmetric, prolate-deformed nucleus is shown in Fig. 1a.
@@
Nuclear shapes, even in ground states, are not fixed. They exhibit zero-point quantum fluctuations involving various collective and nucleonic degrees of freedom (DOF) at different timescales. These fluctuations superimpose on each other in the laboratory frame. In well-deformed nuclei like ${}^{238}$, dominant fluctuations are in the rotational DOF with a timescale of $\tau_{\mathrm{rot}}\sim I/\hbar\sim 10^{3}$$10^{4}$ fm/$c$ (1 fm/$c=3\times 10^{-24}$ seconds) Nakatsukasa et al. (2016), where $I$ denotes the moment of inertia (Fig. 1b). Consequently, measurement processes in spectroscopic methods, lasting orders of magnitude longer than $\tau_{\mathrm{rot}}$, capture a coherent superposition of wavefunctions in all orientations, obscuring instantaneous nuclear shapes. However, these shapes can be extracted precisely by comparing spectroscopic data (Fig. 1c) with model calculations Heyde and Wood (2011); Bender et al. (2003). Traditional electron-nucleus scattering experiments, though faster than $\tau_{\mathrm{rot}}$, probe mostly localized regions of the nucleus, resulting in an orientation-averaged spherical image after accumulating many events, where the impact of nuclear deformation manifests as a broadening of the charge distribution Hofstadter (1956); Bohr and Mottelson (1998).
@@
Fig. 1: Methods for determining the nuclear shape in low and high energies. a, cartoon of a well-deformed prolate-shaped nucleus. b, quantum fluctuations over Euler angles for this nucleus and associated overall timescale. c, quantum mechanical manifestation of the deformation in terms of the first rotational band of ${}^{238}$. d, aligning the two nuclei in headon body-body configuration (top) and tip-tip configuration (bottom). e, collision of two Lorentz-contracted nuclei at high energy and resulting 3D profile of the initially-produced quark-gluon plasma, where arrows indicate the pressure gradients. f, 3D profile of the quark-gluon plasma at the end of the hydrodynamic expansion before freezes out into particles, where arrows indicate the velocities of fluid cells. g, charged particle tracks measured in the detector. The timescales shown are in units of fm/$c$ – the time for light to travel one femtometer. The body-body configuration has large eccentricity $\varepsilon_{2}$ and small gradient $d_{\perp}$, leading to large elliptic flow $v_{2}$ and smaller average transverse momentum $[p_{\mathrm{T}}]$, and verse versa for tip-tip configuration (see text).
@@
New shape-imaging method. To directly observe the nuclei’s global shape, a measurement must 1) be much quicker than $\tau_{\mathrm{rot}}$ and 2) provide access to the many-body nucleon distribution in each nucleus, which contains the frame-independent shape information. High-energy nuclear collisions, an utterly destructive process, remarkably fulfill these criteria. Conducted at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) with center-of-mass energies per nucleon pair ($\sqrt{s_{\mathrm{NN}}}$) reaching up to 5000 GeV, these collisions completely obliterate the nuclei, temporarily forming a quark-gluon plasma (QGP) – a hot, dense matter of interacting quarks and gluons Riordan and Zajc (2006); Gale et al. (2013); Shuryak (2017); Busza et al. (2018). The nuclear shape influences the QGP’s geometry and its collective expansion, imprinting itself on the momentum distribution of the produced particles. In an ironic twist, this effectively realizes Richard Feynman’s analogy of the seemingly impossible task of “figuring out a pocket watch by smashing two together and observing the flying debris,” and the collective response plays a key role.
@@
Our shape-imaging technique focuses on headon (near zero impact parameter) collisions of prolate-deformed nuclei (Fig. 1d-g). The initial configurations lie between two extremes: body-body (top) and tip-tip (bottom) collisions (Fig. 1d). Prior to impact, Lorentz-contraction flattens the ground state nuclei into pancake-like shapes by a factor of $\Gamma=\frac{1}{2}\mbox{$\sqrt{s_{\mathrm{NN}}}$}/m_{0}>100$, where $m_{0}\approx 0.94$ GeV is the nucleon mass (Fig. 1e). The initial impact, lasting $\tau_{\mathrm{expo}}=2R_{0}/\Gamma\lesssim 0.1$ fm/$c$, acts as an exposure time. The shape and size of the overlap region, reflecting the initially-produced QGP, directly mirror those of the colliding nuclei projected on the transverse ($xy$) plane (Fig. 1e). Body-body collisions create a larger, elongated QGP, which undergoes pressure-gradient-driven expansion (indicated by arrows) until about 10 fm/$c$ Busza et al. (2018), resulting in an inverted, asymmetric distribution (Fig. 1f). In contrast, tip-tip collisions form a compact, circular QGP, driving a more rapid but near symmetric expansion (Fig. 1f). In the final stage, the QGP freezes out into thousands of particles, captured as tracks in detectors, whose angular distributions reflect the initial QGP shape (Fig. 1g). This flow-assisted imaging is akin to the Coulomb Explosion Imaging in molecule structure analysis Vager et al. (1989); Kunitski et al. (2015); Endo et al. (2020); Boll et al. (2022); Miller (2022), where the spatial arrangement of atoms, ionized by an X-ray laser or through passage in thin foils, is deduced from their mutual Coulomb-repulsion-driven expansion. However, the expansion duration in high-energy collisions is about $10^{6}$$10^{9}$ times shorter.
@@
Energy evolution of shapes? A pertinent question arises: How do the nuclear shapes observed in high-energy colliders compare to those derived from low-energy experiments? For well-deformed nuclei like ${}^{238}$, we expect them to align at a basic level. However, there are other correlations (collective vibration, clustering, short-range correlations, etc.) that manifest over increasingly faster timescales from 1000 to a few fm/$c$ Nazarewicz (1998). Moreover, high-energy collisions also probe nuclear structure at subnucleonic level, such as quark and gluon correlations, and modifications caused by dense gluon fields Mäntysaari et al. (2023). As a result, the deformations observed at high energy might differ from those at low energy. These factors are subdominant for well-deformed nuclei but could be more discernible for less-deformed ones. This highlights the need to examine nuclear phenomena across energy scales and discover new phenomena.
@@
Observables. In Fig. 1e, the QGP’s initial shape is quantified by the eccentricity, $\varepsilon_{2}=\frac{\left\langle y^{2}\right\rangle-\left\langle x^{2}\right\rangle}{\left\langle y^{2}\right\rangle+\left\langle x^{2}\right\rangle}$, calculated from the nucleon distribution in the $xy$-plane, perpendicular to the beam direction. The hydrodynamic expansion, reacting to $\varepsilon_{2}$, results in particle anisotropy, described as $dN/d\phi~{}\propto~{}1+2v_{2}\cos(2\phi)$ aligned with the impact parameter along the $x$-axis. This phenomenon, known as “elliptic flow” ($v_{2}$llitrault (1992), is depicted in Fig. 1g. Additionally, the QGP’s compactness or density gradient, indicated by the overlap’s inverse area $d_{\perp}=1/\sqrt{\left\langle x^{2}\right\rangle\left\langle y^{2}\right\rangle}$ Schenke et al. (2020a), influences the radial expansion or “radial flow”, captured in the event-wise average transverse momentum ($[p_{\mathrm{T}}]$). A key discovery at RHIC was the QGP’s behavior as a nearly perfect, inviscid fluid Adams et al. (2005); Adcox et al. (2005), effectively transforming initial geometry into final state anisotropies. Hydrodynamic models have confirmed linear response relations: $v_{2}\propto\varepsilon_{2}$ Niemi et al. (2013) and $\delta p_{\mathrm{T}}\propto\delta d_{\perp}$ Boźek and Broniowski (2012), where $\delta p_{\mathrm{T}}=[p_{\mathrm{T}}]-\left\langle[p_{\mathrm{T}}]\right\rangle$ and $\delta d_{\perp}=d_{\perp}-\left\langle d_{\perp}\right\rangle$ denote event-wise deviations from mean values.
@@
In collisions with spherical nuclei, non-zero $\varepsilon_{2}$ and $\delta d_{\perp}$ can be generated by the random fluctuations in the position of nucleons in the overlap region. In non-headon collisions, besides these stochastic elements, the overlap region also has an average elliptical shape. This average shape significantly contributes to $\varepsilon_{2}$, known as reaction plane eccentricity $\varepsilon_{2}^{\mathrm{rp}}$ Alver et al. (2008), but has little effect on $\delta d_{\perp}$.
@@
Prolate deformation in the colliding nuclei further modifies $\varepsilon_{2}$ and $d_{\perp}$. Body-body collisions in this context yield high $\varepsilon_{2}$ and low $d_{\perp}$ values, and vice versa for tip-tip collisions. This leads to enhanced, anti-correlated event-by-event fluctuations in $\varepsilon_{2}$ and $d_{\perp}$ Giacalone (2020a), measurable through observables like $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ that are linearly related to the moments of the initial condition $\left\langle\varepsilon_{2}^{2}\right\rangle$, $\left\langle(\delta d_{\perp})^{2}\right\rangle$, and $\left\langle\varepsilon_{2}^{2}\delta d_{\perp}\right\rangle$. These observables, linked to two- and three-body nucleon distributions in the intrinsic frame (Methods), were found to have a simple parametric dependence on the nuclear shape Jia (2022):
@@
The positive coefficients $a_{n}$ and $b_{n}$ reflect the collision geometry and QGP properties. The $b_{n}$ values are nearly independent of the impact parameter, while $a_{n}$ values are minimized in headon collisions, making such collisions ideal for constraining nuclear shape. Our study represents the first quantitative, simultaneous determination of $\beta_{2}$ and $\gamma$ using all three observables in Eq. (2).
@@
Rigorous calculations of the triaxiality for prolate-like heavy nuclei were limited Hayashi et al. (1984); Sheikh and Hara (1999), with early estimates typically based on simple geometric rotor models, suggesting $\gamma_{\mathrm{U}}=6^{\circ}$$8^{\circ}$ Wu and Cline (1996). A sophisticated energy density functional calculation based on a collective Hamiltonian method (Gogny-5DCH) predicts $\gamma_{\mathrm{U}}\sim 8^{\circ}$ Delaroche et al. (2010). A more recent analysis via a microscopic model suggests a rigid triaxiality with $\gamma_{\mathrm{U}}\sim 15^{\circ}$ Rouoof et al. (2023). However, the origin of triaxiality in these nuclei (rotational or vibrational) remains debated Sharpey-Schafer et al. (2019). Recent advances in the Monte-Carlo Shell Model enabled a more quantitative analysis, indicating a minor rigid triaxial component leading to a small non-zero $\gamma$ for these nuclei, in alignment with spectra data Otsuka et al. (2023). However, triaxiality may change when nuclei are excited, complicating direct spectroscopic verification Otsuka et al. (2023). Thus, measuring ${}^{238}$’s shape could validate our imaging method and directly investigate its ground-state triaxiality.
Fig. 2: Correlation between elliptic flow and radial flow. a, $\left\langle v_{2}^{2}\right\rangle$ versus $\delta p_{\mathrm{T}}/\left\langle[p_{\mathrm{T}}]\right\rangle$ in 0–0.5% most central Au+Au and U+U collisions. b, $\rho_{2}~{}=~{}\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ across centrality, quantifying the strength of $v_{2}$$\delta p_{\mathrm{T}}$ correlation. The elliptic-shaped overlaps in the transverse plane for various centralities are also illustrated.
Our analysis uses U+U data from 2012 and Au+Au data from 2010 and 2011. Each collision produces up to 2000 charged particles in the STAR time-projection chamber (TPC) Anderson et al. (2003), covering the polar angle range $|\theta-90^{\circ}|\lesssim 50^{\circ}$ and full $\phi$ range. The TPC tracks these particles and determines their $p_{\mathrm{T}}$. Collision events are categorized by “centrality”, defined as the percentage of the total inelastic cross-section, with lower percentages indicating more headon collisions. We calculate $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ via established methods Bilandzic et al. (2011) using tracks in $30.2<p_{\mathrm{T}}<$ 3 GeV/$c$. The results incorporate the uncertainties arising from track selection, reconstruction efficiency, background events, and correlations unrelated to flow (Methods).
@@
Results. To directly observe the shape-size correlation depicted in Fig. 1e-f, we analyze the 0–0.5% most central collisions, and correlate $\left\langle v_{2}^{2}\right\rangle$ with event-wise $\delta p_{\mathrm{T}}$ values (Fig. 2a). A pronounced anticorrelation in U+U collisions aligns with the expectation Giacalone (2020a): events with small (large) $\delta p_{\mathrm{T}}$ are enriched with body-body (tip-tip) collisions. This effect is striking, as $\left\langle v_{2}^{2}\right\rangle$ in U+U is twice that of Au+Au at the lowest $\delta p_{\mathrm{T}}$, yet similar at the highest $\delta p_{\mathrm{T}}$.
@@
We quantify this correlation using normalized covariance $\rho_{2}~{}=\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ (Fig. 2b). In Au+Au collisions, $\rho_{2}$ stays relatively constant, with a minor decrease in central regions due to centrality smearing Aad et al. (2023). This smearing can be reduced by averaging over a wider, say 0–5% centrality range Zhou and Jia (2018). In contrast, $\rho_{2}$ in U+U collisions decreases steadily, turning negative at about 7% centrality, reflecting the large prolate deformation of ${}^{238}$. The deformation has the greatest impact in central collisions but also influences other centrality ranges.

### v2

@@ Abstract
Atomic nuclei are self-organized, many-body quantum systems bound by strong nuclear forces within femtometer-scale space. These complex systems manifest a variety of shapes Bohr and Mottelson (1998); Heyde and Wood (2016), traditionally explored using non-invasive spectroscopic techniques at low energies Cline (1986); Yang et al. (2023). However, at these energies, their instantaneous shapes are obscured by long-timescale quantum fluctuations, making direct observation challenging. Here we introduce the “collective flow assisted nuclear shape imaging” method, which images the nuclear global shape by colliding them at ultrarelativistic speeds and analyzing the collective response of outgoing debris. This technique captures a collision-specific snapshot of the spatial matter distribution within the nuclei, which, through the hydrodynamic expansion, imprints patterns on the particle momentum distribution observed in detectors Shuryak (2017); Busza et al. (2018). We benchmark this method in collisions of ground state Uranium-238 nuclei, known for their elongated, axial-symmetric shape. Our findings show a large deformation with a slight deviation from axial symmetry in the nuclear ground state, aligning broadly with previous low-energy experiments. This approach offers a new method for imaging nuclear shapes, enhances our understanding of the initial conditions in high-energy collisions and addresses the important issue of nuclear structure evolution across energy scales.
More than 99.9% of the visible matter in the cosmos resides in the center of atoms – the atomic nuclei composed of nucleons (protons and neutrons). Our knowledge of their global structure primarily comes from spectroscopic or scattering experiments Cline (1986); Yang et al. (2023); Hofstadter (1956) at beam energies below hundreds of MeV per nucleon. These studies show that most nuclei are ellipsoidally deformed, with greater deformation in nuclei distant from magic numbers (2, 8, 20, 28, 50, 82, and 126) Möller et al. (2016). Investigating nuclear shape across the Segrè chart has been an important area of research over many decades and is crucial for topics such as nucleosynthesis Schatz et al. (1998), nuclear fission Schunck and Regnier (2022), and neutrinoless double beta decay ($0\nu\beta\beta$Engel and Menéndez (2017).
In a collective model picture, the ellipsoidal shape of a nucleus with mass number $A$ is defined in the intrinsic (body-fixed) frame, in which its surface $R(\theta,\phi)$ is described by Bohr and Mottelson (1998),
@@
Here $R_{0}\approx 1.2A^{1/3}$ fm represents the nuclear radius. The spherical harmonics in the real basis $Y_{l,m}(\theta,\phi)$, the quadrupole deformation magnitude $\beta_{2}$, and the triaxiality parameter $\gamma$ define the nuclear shape. The $\gamma$ parameter, spanning $0^{\circ}$$60^{\circ}$, controls the ratios of principal radii. Specifically, $\gamma=0^{\circ}$ corresponds to a prolate shape, $\gamma=60^{\circ}$ an oblate shape, and values in between $0^{\circ}<\gamma<60^{\circ}$ to a triaxial shape. Although most nuclei are axially-symmetric (prolate or oblate) or have a fluctuating $\gamma$ value ($\gamma$-soft), the rigid triaxial shape is uncommon Toh et al. (2013). An example of an axial-symmetric, prolate-deformed nucleus is shown in Fig. 1a.
@@
Nuclear shapes, even in ground states, are not fixed. They exhibit zero-point quantum fluctuations involving various collective and nucleonic degrees of freedom (DOF) at different timescales. These fluctuations superimpose on each other in the laboratory frame. In well-deformed nuclei like 238U, dominant fluctuations are in the rotational DOF with a timescale of $\tau_{\mathrm{rot}}\sim I/\hbar\sim 10^{3}$$10^{4}$ fm/$c$ (1 fm/$c=3\times 10^{-24}$ seconds $=$ 3 yoctoseconds) Nakatsukasa et al. (2016), where $I$ denotes the moment of inertia (Fig. 1b). Consequently, measurement processes in spectroscopic methods, lasting orders of magnitude longer than $\tau_{\mathrm{rot}}$, capture a coherent superposition of wavefunctions in all orientations. Their shapes are usually inferred by comparing spectroscopic data (Fig. 1c) with model calculations Heyde and Wood (2011); Bender et al. (2003). Traditional electron-nucleus scattering experiments, although faster than $\tau_{\mathrm{rot}}$, probe mainly localized regions of the nucleus, giving an orientation-averaged spherical image after accumulating many events, in which the impact of deformation manifests as a broadening of the charge distribution Hofstadter (1956); Bohr and Mottelson (1998).
@@
Fig. 1: Methods for determining the nuclear shape in low and high energies. a, Cartoon of a well-deformed prolate-shaped nucleus. b, Quantum fluctuations over Euler angles for this nucleus and associated overall timescale. c, Quantum mechanical manifestation of the deformation in terms of the first rotational band of 238U. d, Aligning the two nuclei in head-on body-body configuration (top) and tip-tip configuration (bottom). e, High-energy collision of two Lorentz-contracted nuclei and resulting 3D profile of the initially-produced quark-gluon plasma, in which the arrows indicate the pressure gradients. f, The 3D profile of the quark-gluon plasma at the end of the hydrodynamic expansion before it freezes out into particles, in which arrows indicate the velocities of fluid cells. g, Charged particle tracks measured in the detector. The timescales shown are in units of fm/$c$ – the time for light to travel 1 femtometer. The body-body configuration has large eccentricity $\varepsilon_{2}$ and small gradient $d_{\perp}$, leading to large elliptic flow $v_{2}$ and smaller average transverse momentum $[p_{\mathrm{T}}]$, and vice versa for tip-tip configuration (see main text).
@@
New shape-imaging method. To directly observe the global shape of the nuclei, a measurement must 1) be much quicker than $\tau_{\mathrm{rot}}$ and 2) provide access to the many-body nucleon distribution in each nucleus. High-energy nuclear collisions, an utterly destructive process, remarkably fulfil these criteria. Conducted at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) with center-of-mass energies per nucleon pair ($\sqrt{s_{\mathrm{NN}}}$) reaching up to 5000 GeV, these collisions completely obliterate the nuclei, temporarily forming a quark-gluon plasma (QGP) – a hot, dense matter of interacting quarks and gluons Shuryak (2017); Busza et al. (2018). The nuclear shape influences the geometry of QGP and its collective expansion, imprinting itself on the momentum distribution of the produced particles. In an ironic twist, this effectively realizes Richard Feynman’s analogy of the seemingly impossible task of “figuring out a pocket watch by smashing two together and observing the flying debris.” The collective response plays a key role.
@@
Our shape-imaging technique focuses on head-on (near zero impact parameter) collisions of prolate-deformed nuclei (Fig. 1d-g). The initial configurations lie between two extremes: body-body (top) and tip-tip (bottom) collisions (Fig. 1d). Before impact, Lorentz-contraction flattens the ground-state nuclei into pancake-like shapes by a factor of $\Gamma=\frac{1}{2}\mbox{$\sqrt{s_{\mathrm{NN}}}$}/m_{0}>100$, where $m_{0}\approx 0.94$ GeV is the nucleon mass (Fig. 1e). The initial impact, lasting $\tau_{\mathrm{expo}}=2R_{0}/\Gamma\lesssim 0.1$ fm/$c$, acts as an exposure time. The shape and size of the overlap region, reflecting the initially produced QGP, directly mirror those of the colliding nuclei projected on the transverse ($xy$) plane (Fig. 1e). Body-body collisions create a larger, elongated QGP, which undergoes pressure-gradient-driven expansion (indicated by arrows) until about 10 fm/$c$ Busza et al. (2018), resulting in an inverted, asymmetric distribution (Fig. 1f). By contrast, tip-tip collisions form a compact, circular QGP, driving a more rapid but symmetric expansion (Fig. 1f). In the final stage, the QGP freezes into thousands of particles, captured as tracks in detectors, whose angular distributions reflect the initial QGP shape (Fig. 1g). This flow-assisted imaging is similar to the Coulomb explosion imaging in molecule structure analysis Vager et al. (1989); Kunitski et al. (2015); Endo et al. (2020); Boll et al. (2022); Miller (2022), in which the spatial arrangement of atoms, ionized by an X-ray laser or through passage in thin foils, is deduced from their mutual Coulomb-repulsion-driven expansion. However, the expansion duration in high-energy collisions is $10^{6}$$10^{9}$ times shorter.
The concept that the dynamics of QGP can be utilized to image the geometrical properties of its initial condition was widely recognized. This understanding has facilitated the determination of the collision impact parameter and fluctuations in nucleon positions Busza et al. (2018), as well as the neutron skin of the colliding nuclei Giacalone et al. (2023), by measuring higher-order harmonic flows. However, our study took a further step to image the shape of the colliding nuclei through their impact on the initial condition.
@@
Energy evolution of shapes? A pertinent question is how the shapes observed in high-energy colliders compare with those derived from low-energy experiments. For well-deformed nuclei like 238U, we expect them to align at a basic level. However, there are other correlations (such as clustering and short-range correlations) that manifest at increasingly faster timescales from 1000 to a few fm/$c$. Moreover, high-energy collisions also probe nuclear structure at sub-nucleonic levels, such as quark and gluon correlations, as well as modifications caused by dense gluon fields Mäntysaari et al. (2023). As a result, the deformations observed at high energy may differ from those at low energy, motivating us to examine nuclear phenomena across energy scales and discover new phenomena.
@@
Observables. In Fig. 1e, the QGP’s initial shape is quantified by the eccentricity, $\varepsilon_{2}=\frac{\left\langle y^{2}\right\rangle-\left\langle x^{2}\right\rangle}{\left\langle y^{2}\right\rangle+\left\langle x^{2}\right\rangle}$, calculated from the nucleon distribution in the $xy$-plane, perpendicular to the beam direction. The hydrodynamic expansion, reacting to $\varepsilon_{2}$, results in particle anisotropy, described as $dN/d\phi~{}\propto~{}1+2v_{2}\cos(2\phi)$ aligned with the impact parameter along the $x$-axis. This phenomenon, known as “elliptic flow” ($v_{2}$llitrault (1992), is shown in Fig. 1g. Moreover, the compactness of the QGP, indicated by the inverse area of the overlap $d_{\perp}\propto 1/\sqrt{\left\langle x^{2}\right\rangle\left\langle y^{2}\right\rangle}$ Schenke et al. (2020a), influences the radial expansion or “radial flow”, captured in the event-wise average transverse momentum ($[p_{\mathrm{T}}]$). A key discovery at RHIC was the behavior of QGP as a nearly perfect, inviscid fluid Adams et al. (2005); Adcox et al. (2005), effectively transforming initial geometry into final state anisotropies. Hydrodynamic models have confirmed linear response relations: $v_{2}\propto\varepsilon_{2}$ Niemi et al. (2013) and $\delta p_{\mathrm{T}}\propto\delta d_{\perp}$ Boźek and Broniowski (2012), where $\delta p_{\mathrm{T}}=[p_{\mathrm{T}}]-\left\langle[p_{\mathrm{T}}]\right\rangle$ and $\delta d_{\perp}=d_{\perp}-\left\langle d_{\perp}\right\rangle$ denote event-wise deviations from mean values.
@@
In head-on collisions with spherical nuclei, non-zero $\varepsilon_{2}$ and $\delta d_{\perp}$ can be generated by the random fluctuations in the position of nucleons in the overlap region. In non-head-on collisions, besides these stochastic elements, the overlap region also has an average elliptical shape. This average shape significantly contributes to $\varepsilon_{2}$, known as reaction plane eccentricity $\varepsilon_{2}^{\mathrm{rp}}$ Alver et al. (2008) but has little effect on the radial quantity $\delta d_{\perp}$.
@@
Prolate deformation further modifies $\varepsilon_{2}$ and $d_{\perp}$. Body-body collisions in this context yield high $\varepsilon_{2}$ and low $d_{\perp}$ values, and vice versa for tip-tip collisions. This leads to enhanced, anti-correlated event-by-event fluctuations in $\varepsilon_{2}$ and $d_{\perp}$ Giacalone (2020a), measurable through observables such as $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ Bozek (2016) that are linearly related to the moments of the initial condition $\left\langle\varepsilon_{2}^{2}\right\rangle$, $\left\langle(\delta d_{\perp})^{2}\right\rangle$, and $\left\langle\varepsilon_{2}^{2}\delta d_{\perp}\right\rangle$. These observables, linked to two- and three-body nucleon distributions in the intrinsic frame (Methods), were found to have a simple parametric dependence on shape parameters Jia (2022):
@@
The positive coefficients $a_{n}$ and $b_{n}$ capture the collision geometry and QGP properties. The $b_{n}$ values are nearly independent of the impact parameter, while $a_{n}$ values are minimized in head-on collisions, making such collisions ideal for constraining nuclear shape. Our study offers the first quantitative and simultaneous determination of $\beta_{2}$ and $\gamma$ using all three observables in Eq. (2).
Nuclear shapes from low energy data. Our measurements use data from high-energy 238U+238U and 197Au+197Au collisions. These species have contrasting shapes: mildly oblate 197Au (close to magic numbers with $Z$=79 protons and $N$=118 neutrons) and highly prolate 238U (an open shell nucleus with 92 protons and 146 neutrons). This comparison helps us to deduce the shape of 238U. A state-of-the-art beyond mean-field model, which reproduces the bulk of experimental data on 197Au, predicts deformation values of $\beta_{2\mathrm{Au}}\approx 0.12$–0.14 and $\gamma_{\mathrm{Au}}\approx 43^{\circ}$ Bally et al. (2023). The deformation of 238U, inferred from measured transition rates within rotational spectra, is estimated at $\beta_{2\mathrm{U}}=0.287\pm 0.007$ Pritychenko et al. (2016).
Experimental estimates of the Uranium triaxiality have been derived from energy levels and transition data under a rigid-rotor assumption, suggesting $\gamma_{\mathrm{U}}=6^{\circ}$$8^{\circ}$ Wu and Cline (1996). One important issue concerns the softness of $\gamma$: whether the nuclei have rigid triaxial shape or fluctuations of $\gamma$ around its mean value Sharpey-Schafer et al. (2019). This issue is complicated by possible changes of $\gamma$ when nuclei are excited Otsuka et al. (2023). Our three-body observable $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in Eq. (2) is sensitive only to the mean of the triaxiality, not its fluctuations Dimri et al. (2023). Nevertheless, measuring $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$ could validate our imaging method and investigate its ground-state triaxiality.
Fig. 2: Correlation between elliptic flow and radial flow. a, $\left\langle v_{2}^{2}\right\rangle$ versus $\delta p_{\mathrm{T}}/\left\langle[p_{\mathrm{T}}]\right\rangle$ in 0–0.5% most central Au+Au and U+U collisions. b, $\rho_{2}~{}=~{}\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ across centrality, quantifying the strength of $v_{2}$$\delta p_{\mathrm{T}}$ correlation. The elliptic-shaped overlaps in the transverse plane for various centralities are also shown.
@@
Results. To directly observe the shape-size correlation depicted in Fig. 1e–f, we analyze the 0–0.5% most central collisions and correlate $\left\langle v_{2}^{2}\right\rangle$ with event-wise $\delta p_{\mathrm{T}}$ values (Fig. 2a). A pronounced anticorrelation in U+U collisions aligns with the expectation Giacalone (2020a): events with small (large) $\delta p_{\mathrm{T}}$ are enriched with body-body (tip-tip) collisions. This effect is striking, as $\left\langle v_{2}^{2}\right\rangle$ in U+U is twice that of Au+Au at the lowest $\delta p_{\mathrm{T}}$, yet similar at the highest $\delta p_{\mathrm{T}}$.
@@
We quantify this correlation using normalized covariance $\rho_{2}~{}=\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ (Fig. 2b). In Au+Au collisions, $\rho_{2}$ is relatively constant, with a minor decrease in the central region due to centrality smearing Aad et al. (2023). This smearing can be reduced by averaging over a wider, say 0–5% centrality range Zhou and Jia (2018). By contrast, $\rho_{2}$ in U+U collisions decreases steadily, turning negative at about 7% centrality, reflecting the large prolate deformation of 238U. The deformation has the greatest impact in central collisions but also influences other centrality ranges.